OpenStudy (alex6799):

Find the values of the 30th and 90th percentiles of the data. Please show your work. 129, 113, 200, 100, 105, 132, 100, 176, 146, 152

OpenStudy (alex6799):

100, 100, 105, 113, 129, 132, 146, 152, 176, 200

OpenStudy (freckles):

ok it looks like you have ordered the data from least to greatest

OpenStudy (freckles):

and that is the first step

OpenStudy (alex6799):

ok good, but i have no idea where to go after that

OpenStudy (freckles):

the second step is trying to figure out what number to use for the 30th percentile which number in the data gives us 30 percent of the elements of the data below it

OpenStudy (freckles):

you have 10 elements

OpenStudy (freckles):

you want to multiply this number by 30%

OpenStudy (freckles):

this will tell us which number in the data is going to have 30% of the elements below it

OpenStudy (alex6799):

10*30%=3

OpenStudy (alex6799):

so 105?

OpenStudy (freckles):

100, 100, 105, 113, 129, 132, 146, 152, 176, 200 so 105 is the third number the only thing is there are two elements below it and 2 elements of a 10 piece data set is not 30% of the data

OpenStudy (freckles):

so we find the average of 105 and the next number in the data and use that as our 30 percentile number

OpenStudy (alex6799):

so 105+113=218 218/2=109

OpenStudy (freckles):

ok that looks great

OpenStudy (alex6799):

YAY!.....now what?

OpenStudy (freckles):

do similar thing for finding 90th percentile number

OpenStudy (freckles):

the number you choose you want 90% of the elements laying below it (instead of 30% like we did with the 30th percentile number)

OpenStudy (alex6799):

so 10*90%=9 #9=176, #10=200 176+200=376 376/2=188

OpenStudy (freckles):

sorry the notification thingy didn't work

OpenStudy (alex6799):

its fine

OpenStudy (freckles):

that looks great to me

OpenStudy (alex6799):

ok thanks!

OpenStudy (freckles):

http://www.dummies.com/how-to/content/how-to-calculate-percentiles-in-statistics.html another very followable example here if you want to read

OpenStudy (freckles):

anyways great job

OpenStudy (alex6799):

thanks, thank you for explaining it to me